OpenStudy (anonymous):

two functions f and g are defined by f: x maps onto 1/ x+1 and g:x maps onto x/x-2. find values of x for which f(x)=10g(x)

OpenStudy (anonymous):

Basically you solve for \(x\): \[\frac{1}{x+1}=\frac{10x}{x-2}\]

OpenStudy (anonymous):

thanx i'll get it from here :)